Given IH = $1,500, IS = $230, and p = 7%, what is the value of E(I)?

$1,411.10

$1,388.90

$1,234

$318.90

none of the above
Answer #1

Given IH=$1500

IS=$230

Probability of being sick=p=7%

Expected value of I=E(I)=p*IS+(1-p)*IH=0.07*230+(1-0.07)*1500=$1411.10

Correct option is

$1411.10